Please help me solve this problem.
In a certain linear pair, one angle measures twice as much as the other, what is the measure of each angle?

Found 2 solutions by malakumar_kos@yahoo.com, maths0503:
Answer by malakumar_kos@yahoo.com(315)


In a certain linear pair, one angle measures twice as much as the other, what is the measure of each angle?

Let the linear pair of angles be x & y
given one angle is two times the other
i.e, x = 2y

linear pair is x+y = 180
2y+y = 180
3y = 180
y = 180/3 = 60
x = 2y = 2*60 = 120
The required angles are 60 and 120


Answer by maths0503(3)

x+2x=180
x=60